10 March 2014

The return of clouds

Eureka 2014: Daily Report for 10 March 2014 The return of clouds Today, the weather at Eureka was cloudy, with temperatures around -41°C. PEARL continues to enjoy the temperature inversion advantage, and was -21°C.

This morning, Paul found MAESTRO lying on the Ridge Lab roof. It must have fallen off the tracker sometime overnight. Winds were not excessive, but Paul speculates that the locking mechanism could have loosened due to repeated temperature changes. Paul and Sophie re-secured MAESTRO to its tracker. The instrument continued to communicate with its computer and take spectra. Paul will re-align the instrument when sunlight returns; hopefully this will be tomorrow morning.

In the PEARL Ridge Lab's IR lab, clouds limited solar measurements to the first 45 minutes at the lab. During that window, Dan took 10 measurements with the Bruker remotely with his computer, and 7 measurements with PARIS. Joseph ran the suntracker remotely from his computer. Once measurements were no longer possible, Joseph worked on the Standard Operating Procedures document for remote suntracker operation. Dan marked term tests remotely and worked on new public outreach articles.

Over in the UV-Vis Lab, Sophie had an issue with the PEARL-GBS shutter. She determined that the shutter and wires were ok, but the code disabled the "close shutter" command. Sophie is working to figure out how to enable it again or change the code.

Next summer, Xiaoyi and Sophie would like to move the suntracker and the PEARL-GBS telescope to the penthouse. They will keep the instrument itself in the UV-Vis lab's mezzanine, as it is now. In preparation, Sophie and Paul looked at the building plans to get a rough idea of how long the optical fibre bundle is needed to connect the two locations. They then went outside and measured the distance between the two locations.

Down at 0PAL, Chris Perro, working remotely from Dalhousie University, performed overlap tests at the beginning of the day when the sky was clear. As conditions turned cloudy, CRL was shifted to routine measurements. This was the 5th day of continuous lidar operation.

Today's ACE ozonesonde flight was launched at 6:15 PM (local time) and reached an altitude of 26 708 m (16.5 hPa).

In the evening, the team watched the traditional ACE Campaign movie, The Thing.

- Dan
Back to Daily Reports
Back to Home